# Stochastic surfing turbulent vorticity

### v1

@@ Abstract
The chaotic dynamics of small-scale vorticity plays a key role in understanding and controlling turbulence, with direct implications for energy transfer, mixing, and coherent structure evolution. However, measuring or controlling its dynamics remains a major conceptual and experimental challenge due to its transient and chaotic nature. Here we use a combination of experiments, theory and simulations to show that small magnetic particles of different densities, exploring flow regions of distinct vorticity statistics, can act as effective probes for measuring and forcing turbulence at its smallest scale. The interplay between the magnetic torque, from an externally controllable magnetic field, and hydrodynamic stresses, from small-scale turbulent vorticity, uncovers an extremely rich phenomenology. Notably, we present the first observation of stochastic resonance for particles in turbulence, where turbulent fluctuations, remarkably acting as an effective noise, enhance the particle rotational response to external forcing. We identify a pronounced resonant peak in the particle rotational phase lag when the applied rotating magnetic field matches the characteristic intensity of small-scale turbulent vortices. Furthermore, we reveal a novel symmetry-breaking mechanism: an oscillating magnetic field with zero-mean angular velocity can counterintuitively induce a net particle rotation in turbulence with zero-mean vorticity, as turbulent fluctuations aid the particle in “surfing” the magnetic field. Leveraging this stochastic resonance, our findings pave the way to the development of techniques for flexibly manipulating particle dynamics in complex flows. Moreover, the discovered mechanism introduces the possibility to develop a novel magnetic resonance-based approach for measuring turbulent vorticity, where particles acting as probes emit a detectable magnetic field that can be used to characterize turbulence even under optically inaccessible conditions.
@@
Figure 1: Rotational dynamics of magnetic particles in a rotating magnetic field: “phase-locked” versus “back-and-forth” regimes. a, Magnetic particles in turbulence under a rotating magnetic field: The experimental setup consists of a turbulence generator, a magnetic field generator, and magnetic particles. A Von Kármán-type turbulent flow is generated in an octagonal water-filled container, driven by two counter-rotating bladed disks. A uniform rotating magnetic field $\bm{H}=H\bm{h}$, with angular frequency $\omega_{H}$, is generated using a system of two pairs of perpendicular Helmholtz coils. b, Magnetic particles consist of a Styrofoam core coated with magnetic paint. Surface patterns enable tracking rotational motion, as illustrated in the stroboscopic time-lapse trajectories (c-d). e, Photograph of the setup. f-i, In a quiescent fluid, particles exhibit two rotational regimes: “phase-locked” (low $\omega_{H}$) and “back-and-forth” (high $\omega_{H}$). f (Supplementary Movie 1) and h (Supplementary Movie 2) show rotational trajectories within one rotation period, viewed in a plane perpendicular to the rotation plane of the magnetic field. The particle initial preferred magnetization direction $\bm{n}_{0}$ (dashed arrows) and instantaneous orientation $\bm{n}$ (solid arrows) are marked over time. During rotation within a quiescent fluid, $\bm{n}$ and $\bm{h}$ remain within the same plane [8]. The corresponding normalized particle angular velocity, $\omega_{p,i}/\omega_{H}$ ($i=x,y,z$), is shown in g and i. Detailed information about the experiments and simulations can be found in Methods.
@@
Figure 2: Experimental validation of the theoretical model for particle rotational dynamics with turbulence. a, Schematic of a magnetic particle in turbulence under a rotating magnetic field. The particle with preferred magnetization direction $\bm{n}$ is subjected to a magnetic field $\bm{h}$ in the $xoy$-plane, rotating around the $z$-axis with frequency $\omega_{H}$, and experiences turbulent vorticity $\bm{\omega}_{f}$ in homogeneous isotropic turbulence. The phase lag $\beta$ (defined as the angle between the projection of $\bm{n}$, i.e., $\bm{n}_{xoy}$, and $\bm{h}$) and the rotation angle $\psi$ (angle of $\bm{n}_{xoy}$ relative to the $x$-axis) describe the particle dynamics. The dynamical evolution of $\beta$ is governed by the interplay of the magnetic field rotation drag, magnetic torque, and turbulence torque. b-d, Comparison of the probability density distribution function (PDF) of particle angular velocity in experimental (circles) and numerical (squares) studies. The experimental results (circles) are shown for varying magnetic field strengths with fixed turbulence intensity. b weak (1.2mT, Supplementary Movie 3), c intermediate (1.4mT, Supplementary Movie 4), and d strong (1.6mT, Supplementary Movie 5), with constant rotational frequency of the magnetic field (parameter settings can be found in the Methods and Supplementary Information).
@@
Figure 4: Stochastic resonance. a, Phase diagram of stochastic resonance. The color represents the normalized time-averaged phase lag derivative, $\langle\dot{\beta}\rangle/\omega_{\eta}$. The vertical axis, $\omega_{\eta}/\omega_{a}$, quantifies the relative intensity of turbulent fluctuations (noise) to the magnetic field strength. The horizontal axis, $\omega_{H}/\omega_{a}$, distinguishes subcritical, critical ($\omega_{H}/\omega_{a}=1/2$, marked by the black dot-dashed line) and supercritical regimes. For a fixed $\omega_{H}/\omega_{a}$ in the subcritical (supercritical) range, $\langle\dot{\beta}\rangle/\omega_{\eta}$ exhibits a non-monotonic (monotonic) dependence on noise. A pronounced resonant peak appears at $\omega_{\eta}/\omega_{a}\approx 1$ when $\omega_{H}/\omega_{a}<1/2$ (b), with a perfect collapse in the linear regime when normalized by $\omega_{H}/\omega_{a}$ (b, inset). c, Effective potential well framework: the magnetic field contributes a cosine potential modulated by a linear term, and turbulence acts as a stochastic noise, see Eq. (2). When $\omega_{H}/\omega_{a}<1/2$ (d), for weak noise, the phase lag remains near the local minimum of the potential well. As noise increases, the particle phase lag exhibits unidirectional escape, signaling stochastic resonance. At even higher noise levels, escape becomes bidirectional, leading to a subsequent decline in $\langle\dot{\beta}\rangle$. For $\omega_{H}/\omega_{a}\ll 1/2$, the potential well is sufficiently deep to strongly localize the particle phase lag (e). At the critical threshold $\omega_{H}/\omega_{a}=1/2$, the potential well vanishes, making the phase lag sensitive to small perturbations (f). In the supercritical range ($\omega_{H}/\omega_{a}\gg 1/2$), the potential well inherently favors persistent phase slipping, regardless of noise intensity (g).
@@
Previous studies have extensively explored magnetic particles in quiescent fluids under rotating magnetic fields, revealing rich phenomena such as self-assembly into chain-like structures [12, 13], synchronization-selected structures [14], and more complex arrays [15], enhanced mixing [12, 16, 13], and phase locking, where particles follow the rotating magnetic field with a constant phase lag with respect to the magnetic field when the rotating magnetic angular velocity $\omega_{H}$ is below a threshold $\omega_{cr}$ [17]. Investigations have also extended to self-propelling active magnetic particles (e.g., magnetotactic bacteria) [18, 19], magnetic Janus particles [14, 20], and particles with different aspect ratios [21, 22].
@@
Here, we explore this question by investigating the rotational dynamics of magnetic particles in turbulence under a rotating magnetic field. We show that the effect of turbulence (vorticity) acts as an effective noise (Supplementary Information) and reveal the first signature of the emergence of stochastic resonance (SR)—a counterintuitive phenomenon where internal fluctuations non-linearly cooperate with external forcing to amplify the system response [23, 24, 25, 26, 27, 28, 29, 30, 31]—of particles in turbulence, which can be used as the mechanism of magnetic particles acting as effective vorticity probes for a novel type of turbulence “microscope”. While the concept of SR, initially observed in bistable systems subjected to noise and weak periodic signals, has since been extended to a wide range of physical and biological systems, as well as signal processing and computer science [32, 33, 34, 35, 36, 37], its role in turbulence-driven rotational dynamics remains mostly unexplored.
To identify potential SR signatures, we set out to leverage turbulent vorticity fluctuations as a source of noise and examine the response of the magnetic particle dynamics to the rotating magnetic field, by means of experiments, simulations, and theoretical modeling. Our experimental platform (Fig. 1a and Fig. 1e) involves a dilute collection of approximately spherical light magnetic particles (produced by Styrofoam core coated with magnetic paint, Fig. 1b, with a mean density of 20% with respect to water and with a mean diameter of 0.7 mm, see Methods for details) which are immersed in a Von Kármán-type turbulent flow (producing homogeneous and isotropic turbulence (HIT) in the center region of the cell [38, 39, 40, 41]) in an octagonal water-filled chamber. The flow is generated by two counter-rotating bladed disks and the chamber is placed in a system of Helmholtz coils, which is programmed to generate a uniform magnetic field in a restricted measurement volume, $\bm{H}=H\bm{h}$, rotating in the $xoy$-plane with angular velocity vector $\bm{\omega}_{H}=\omega_{H}\hat{\bm{z}}$ where $\bm{\hat{z}}$ is the unit vector along the $z$-axis (Methods and Supplementary Information). As the induced magnetic dipole, $\bm{m}$, of a particle tends to align with the magnetic field direction, $\bm{h}$, the particle experiences a magnetic torque balanced by the rotational drag and spins around an axis in the particle body frame parallel to $\hat{\bm{z}}$ along its trajectory (as shown by typical stroboscopic time-lapse trajectories in Fig. 1c-d). The rotational motion is measured by tracking particle surface patterns (Methods).
@@
The critical frequency is $\omega_{cr}=\omega_{a}/2$ (obtained in a noiseless system, i.e., $\bm{\omega}_{f}=\bm{0}$, details in Supplementary Information). For $\omega_{H}\leq\omega_{cr}$, the particle is synchronized with the rotating magnetic field (“phase-locked”), i.e. $\omega_{p,z}=\omega_{H}$. Otherwise, when $\omega_{H}>\omega_{cr}$, the particle undergoes back-and-forth motion (“back-and-forth”) when the maximum magnetic torque cannot balance the response drag torque: the phase lag angle $\beta$ between $\bm{n}_{xoy}$ (the projection of $\bm{n}$ on $xoy$-plane) and $\bm{h}$ changes periodically over time. These two regimes can be observed experimentally (Fig. 1f-i). In the “phase-locked” regime (Supplementary Movie 1), the particle rotates synchronously with the magnetic field, as evidenced by the normalized angular velocity, $\omega_{p,z}/\omega_{H}\approx 1$, in Fig. 1g. In the “back-and-forth” regime (Supplementary Movie 2), the particle oscillates periodically (Fig. 1i).
As illustrated in the schematic diagram of Fig. 2a, the dynamics of the particle preferred magnetization direction, $\bm{n}$, is fundamentally characterized by the phase lag angle, $\beta$. The governing equations, Eqs. (1), can be simplified and explicitly written as (derivations in Methods)
@@
Geometrically, we have $\beta=\omega_{H}t-\psi$. The phase lag angle $\beta$ represents the particle rotational displacement in the co-rotating frame of the magnetic field, while $\psi$ is the rotation angle of the projection $\bm{n}$ on the $xoy$-plane with respect to the $x$-axis, which describes the particle rotational displacement in the laboratory frame. The turbulent vorticity signal, $\bm{\omega}_{f}$, represents the vorticity experienced by the particle center along its trajectory. This simplified representation, Eq. (2), separates different contributions and clearly shows that the rotational motion of the particle results from the interplay between the magnetic field rotational drag, the magnetic torque, and the turbulence drag torque.
@@
Now we wonder how this picture will change when turbulence is taken into consideration. The turbulence induces fluctuations on the small particles of a typical magnitude $\omega_{\eta}=1/\tau_{\eta}$, the Kolmogorov characteristic frequency. Experimentally, we turn on the turbulence and vary the intensity of the magnetic field. By statistically analyzing the probability density function (PDF) of the particle angular velocity, a dominant peak is observed in the $z$-direction at approximately $\omega_{p,z}\approx\omega_{H}$ (Fig. 2b-d, purple circles), while the angular velocities in the $x$ and $y$ directions are similar to each other but significantly different from that in the $z$-direction (Fig. 2b-d, yellow and green circles).
To understand the influence of turbulence on the particle dynamics further, we carry out a systematic numerical investigation. To model the motion of light particles in homogeneous isotropic turbulence, we first perform direct numerical simulations using a pseudo-spectral method [43, 44, 11] to obtain $\bm{\omega}_{f}$ (see Methods for details). The resulting vorticity data is then incorporated into the theoretical model, Eqs. (1), which is integrated using a third-order Runge-Kutta method.
The simulated particle rotation is validated against the experimental results (Fig. 2b-d, Supplementary Movies 3-5, respectively). Notably, the theoretical model (squares) accurately captures the essential statistical physics (peak positions) of the particle stochastic rotational dynamics in the experiments (circles). The slight discrepancies in the PDF shapes arise from polydispersity in particle shape and magnetic properties in experiments which influences how individual particles respond to an external magnetic field, whereas the simulations assume idealized spherical particles with the same magnetic characteristics for each particle (Methods and Supplementary Information). Further comparisons between simulated and experimental results all demonstrate good qualitative agreement (Supplementary Information, Supplementary Movies 9 and 10).
@@
With the theoretical model capturing the core dynamics of this complex system, and with simulations allowing exploration of a broader parameter space and finer details of particle rotation, we systematically investigate how turbulence-induced vorticity fluctuations and the deterministic forcing of the applied magnetic field shape the particle rotational dynamics.
[…]
@@
To determine whether SR occurs, we focus on the phase lag angle and its behavior across different parameter settings. We summarize the time-averaged value of $\langle\dot{\beta}\rangle/\omega_{\eta}$ in one phase diagram (Fig. 4a), with $\omega_{\eta}/\omega_{a}$ (vertical axis) quantifying the relative turbulent fluctuation intensity (noise) to the applied magnetic field strength and $\omega_{H}/\omega_{a}$ (horizontal axis) distinguishing noiseless subcritical, critical ($\omega_{H}/\omega_{a}=1/2$, marked by the black dot-dashed line), and supercritical states.
@@
For a fixed $\omega_{H}/\omega_{a}$ in the subcritical regime, the particle rotation in the co-rotating frame $\langle\dot{\beta}\rangle/\omega_{\eta}$ exhibits a non-monotonic dependence on noise intensity (Fig. 4b), with a pronounced resonant peak at $\omega_{\eta}/\omega_{a}\approx 1$ when $\omega_{H}/\omega_{a}<1/2$, indicating that resonance arises at the transition from the “phase-locked” regime to the “turbulent-dynamics” regime. This is a clear suggestion that the observed resonance behavior results from the cooperation between turbulent fluctuations and magnetic forcing. The resonance mechanism can be interpreted via an effective potential well framework (Fig. 4c), where the magnetic field induces a cosine-type potential (i.e., $-\frac{\omega_{a}}{4}\cos{2\beta}$) modulated by a linear term (i.e., $-\omega_{H}\beta$), while turbulence acts as a stochastic noise.
@@
For $\omega_{H}/\omega_{a}<1/2$ (Fig. 4d), the phase lag angle is initially trapped in the potential minimum at weak noise levels (small $\omega_{\eta}/\omega_{a}$), exhibiting only mild fluctuations locally. As the turbulence-induced fluctuations become comparable to the magnetic field strength ($\omega_{\eta}/\omega_{a}\approx 1$), preferential unidirectional escape occurs. This clearly identifies the observed resonant behavior as an example of stochastic resonance: a cooperative interplay between stochastic forcing and the external deterministic forcing. This phenomenon can be thought of as the particle “surfing” the rotating magnetic field, aided by the turbulent vorticity. At even higher noise levels, randomization overrides the cooperative effect, causing bidirectional escape and a decline in $\langle\dot{\beta}\rangle$.
@@
At the critical threshold $\omega_{H}/\omega_{a}=1/2$, the potential well vanishes, rendering $\beta$ critically sensitive to any perturbations (Fig. 4f). In the supercritical range ($\omega_{H}/\omega_{a}\gg 1/2$), the potential well inherently favors persistent phase slipping, regardless of noise intensity (Fig. 4g), such that the SR mechanism vanishes in this regime.

### v2

# Stochastic resonance of rotating particles in turbulence

@@ Abstract
The chaotic dynamics of small-scale vorticity plays a key role in understanding and controlling turbulence, with direct implications for energy transfer, mixing, and coherent structure evolution. However, measuring or controlling its dynamics remains a major conceptual and experimental challenge due to its transient and chaotic nature. Here we use a combination of experiments, theory and simulations to show that small magnetic particles of different densities, exploring flow regions of distinct vorticity statistics, can act as effective probes for measuring and forcing turbulence at its smallest scale. The interplay between the magnetic torque, from an externally controllable magnetic field, and hydrodynamic stresses, from small-scale turbulent vorticity, uncovers an extremely rich phenomenology. Notably, we present the first experimental and numerical observation of stochastic resonance for particles in turbulence, where turbulent fluctuations, remarkably acting as an effective noise, enhance the particle rotational response to external forcing. We identify a pronounced resonant peak in the particle rotational phase lag when the applied rotating magnetic field matches the characteristic intensity of small-scale turbulent vortices. Furthermore, we reveal a novel symmetry-breaking mechanism: an oscillating magnetic field with zero-mean angular velocity can counterintuitively induce a net particle rotation in turbulence with zero-mean vorticity, as turbulent fluctuations aid the particle in “surfing” the magnetic field. Leveraging this stochastic resonance, our findings pave the way to the development of techniques for flexibly manipulating particle dynamics in complex flows. Moreover, the discovered mechanism introduces the possibility to develop a novel magnetic resonance-based approach for measuring turbulent vorticity, where particles acting as probes emit a detectable magnetic field that can be used to characterize turbulence even under optically inaccessible conditions.
@@
Figure 1: Rotational dynamics of magnetic particles in a rotating magnetic field: “phase-locked” versus “back-and-forth” regimes. a, Magnetic particles in turbulence under a rotating magnetic field: The experimental setup consists of a turbulence generator, a magnetic field generator, and magnetic particles. A Von Kármán-type turbulent flow is generated in an octagonal water-filled container with an internal diameter of $2R=150$ mm and a height of 220 mm, driven by two counter-rotating bladed disks. A uniform rotating magnetic field $\bm{H}=H\bm{h}$, with angular frequency $\omega_{H}$, is generated using a system of two pairs of perpendicular Helmholtz coils. b, Magnetic particles consist of a Styrofoam core coated with magnetic paint. Surface patterns enable tracking rotational motion, as illustrated in the stroboscopic time-lapse trajectories (c-d). e, Photograph of the setup. f-i, In a quiescent fluid, particles exhibit two rotational regimes: “phase-locked” (low $\omega_{H}$) and “back-and-forth” (high $\omega_{H}$). f (Supplementary Movie 1) and h (Supplementary Movie 2) show rotational trajectories within one rotation period, viewed in a plane perpendicular to the rotation plane of the magnetic field. The particle initial preferred magnetization direction $\bm{n}_{0}$ (dashed arrows) and instantaneous orientation $\bm{n}$ (solid arrows) are marked over time. During rotation within a quiescent fluid, $\bm{n}$ and $\bm{h}$ remain within the same plane [8]. The corresponding normalized particle angular velocity, $\omega_{p,i}/\omega_{H}$ ($i=x,y,z$), is shown in g and i. Detailed information about the experiments and simulations can be found in Methods.
@@
Figure 2: Experimental validation of the theoretical model for particle rotational dynamics with turbulence. a, Schematic of a magnetic particle in turbulence under a rotating magnetic field. The particle with preferred magnetization direction $\bm{n}$ is subjected to a magnetic field $\bm{h}$ in the xoyxoy-plane, rotating around the $z$-axis with frequency $\omega_{H}$, and experiences turbulent vorticity $\bm{\omega}_{f}$ in homogeneous isotropic turbulence. The phase lag $\beta$ (defined as the angle between the projection of $\bm{n}$, i.e., $\bm{n}_{xoy}$, and $\bm{h}$) and the rotation angle $\psi$ (angle of $\bm{n}_{xoy}$ relative to the $x$-axis) describe the particle dynamics. The dynamical evolution of $\beta$ is governed by the interplay of the magnetic field rotation drag, magnetic torque, and turbulence torque. b-d, Comparison of the probability density distribution function (PDF) of particle angular velocity in experimental (circles) and numerical (squares) studies. The experimental results (circles) are shown for varying magnetic field strengths with fixed turbulence intensity. b weak (1.2mT, Supplementary Movie 3), c intermediate (1.4mT, Supplementary Movie 4), and d strong (1.6mT, Supplementary Movie 5), with constant rotational frequency of the magnetic field (parameter settings can be found in the Methods and Supplementary Information).
@@
Figure 4: Stochastic resonance. a, Phase diagram of stochastic resonance. The color represents the normalized time-averaged phase lag derivative, $\langle\dot{\beta}\rangle/\omega_{\eta}$. The vertical axis, $\omega_{\eta}/\omega_{a}$, quantifies the relative intensity of turbulent fluctuations (noise) to the magnetic field strength. The horizontal axis, $\omega_{H}/\omega_{a}$, distinguishes subcritical, critical ($\omega_{H}/\omega_{a}=1/2$, marked by the black dot-dashed line) and supercritical regimes. For a fixed $\omega_{H}/\omega_{a}$ in the subcritical (supercritical) range, $\langle\dot{\beta}\rangle/\omega_{\eta}$ exhibits a non-monotonic (monotonic) dependence on noise. A pronounced resonant peak appears at $\omega_{\eta}/\omega_{a}\approx 1$ when $\omega_{H}/\omega_{a}<1/2$ (b), with a perfect collapse in the linear regime when normalized by $\omega_{H}/\omega_{a}$ (b, inset). The diamond symbols in panel a denote the parameter settings for experiments shown in panel c. c, Experimental evidence of stochastic resonance (diamonds) at $\omega_{H}/\omega_{a}=0.12\pm 0.01$. For comparison, simulation results at a similar value of $\omega_{H}/\omega_{a}$ (circles) are also shown. The error bars correspond to the standard deviation calculated from 20 independent measurements at the same parameter value. d, Effective potential well framework: the magnetic field contributes a cosine potential modulated by a linear term, and turbulence acts as a stochastic noise, see Eq. (2). When $\omega_{H}/\omega_{a}<1/2$ (e), for weak noise, the phase lag remains near the local minimum of the potential well. As noise increases, the particle phase lag exhibits unidirectional escape, signaling stochastic resonance. At even higher noise levels, escape becomes bidirectional, leading to a subsequent decline in $\langle\dot{\beta}\rangle$. For $\omega_{H}/\omega_{a}\ll 1/2$, the potential well is sufficiently deep to strongly localize the particle phase lag (f). At the critical threshold $\omega_{H}/\omega_{a}=1/2$, the potential well vanishes, making the phase lag sensitive to small perturbations (g). In the supercritical range ($\omega_{H}/\omega_{a}\gg 1/2$), the potential well inherently favors persistent phase slipping, regardless of noise intensity (h).
@@
Previous studies have extensively explored magnetic particles in quiescent fluids under rotating magnetic fields, revealing rich phenomena such as self-assembly into chain-like structures [12, 13], synchronization-selected structures [14], and more complex arrays [15], enhanced mixing [12, 16, 13], turbulence design [17, 18, 19] and phase locking, where particles follow the rotating magnetic field with a constant phase lag with respect to the magnetic field when the rotating magnetic angular velocity $\omega_{H}$ is below a threshold $\omega_{cr}$ [20]. Investigations have also extended to self-propelling active magnetic particles (e.g., magnetotactic bacteria) [21, 22], magnetic Janus particles [14, 23], particles with different aspect ratios [24, 25], colloidal spinners and active swimmers [26, 27, 28].
@@
Here, we explore this question by investigating the rotational dynamics of magnetic particles in turbulence under a rotating magnetic field.
While deterministic synchronized and oscillation (back-and-forth) rotational regimes have been observed in zero-noise systems (e.g., compasses or rolling particles in quiescent fluids [29, 30, 31]), our work extends this understanding into the complex environment of fully developed turbulence. We systematically explore the complex interplay between deterministic external magnetic forcing and turbulent hydrodynamic torques, revealing a significantly richer phase space of particle rotational dynamics. We show that the effect of turbulence (vorticity) acts as an effective noise (Supplementary Information) and reveal the first signature of the emergence of stochastic resonance (SR)—a counterintuitive phenomenon where internal fluctuations non-linearly cooperate with external forcing to amplify the system response [32, 33, 34, 35, 36, 37, 38, 39, 40]—of particles in turbulence, which can be used as the mechanism of magnetic particles acting as effective vorticity probes for a novel type of turbulence “microscope”. While the concept of SR, initially observed in bistable systems subjected to noise and weak periodic signals, has since been extended to a wide range of physical and biological systems, as well as signal processing and computer science [41, 42, 43, 44, 45, 46], its role in turbulence-driven rotational dynamics remains mostly unexplored.
To identify potential SR signatures, we set out to leverage turbulent vorticity fluctuations as a source of noise and examine the response of the magnetic particle dynamics to the rotating magnetic field, by means of experiments, simulations, and theoretical modeling. Our experimental platform (Fig. 1a and Fig. 1e) involves a dilute collection of approximately spherical light magnetic particles (produced by Styrofoam core coated with magnetic paint, Fig. 1b, with a mean density of 20% with respect to water and with a mean diameter of 0.7 mm, see Methods for details) which are immersed in a Von Kármán-type turbulent flow (producing homogeneous and isotropic turbulence (HIT) in the center region of the cell [47, 48, 49, 50]) in an octagonal water-filled chamber. The flow is generated by two counter-rotating bladed disks and the chamber is placed in a system of Helmholtz coils, which is programmed to generate a uniform magnetic field in a restricted measurement volume, $\bm{H}=H\bm{h}$, rotating in the xoyxoy-plane with angular velocity vector $\bm{\omega}_{H}=\omega_{H}\hat{\bm{z}}$ where $\bm{\hat{z}}$ is the unit vector along the $z$-axis (Methods and Supplementary Information). As the induced magnetic dipole, $\bm{m}$, of a particle tends to align with the magnetic field direction, $\bm{h}$, the particle experiences a magnetic torque balanced by the rotational drag and spins around an axis in the particle body frame parallel to $\hat{\bm{z}}$ along its trajectory (as shown by typical stroboscopic time-lapse trajectories in Fig. 1c-d). The rotational motion is measured by tracking particle surface patterns (Methods).
@@
In a noiseless system, i.e., $\bm{\omega}_{f}=\bm{0}$, the critical frequency is $\omega_{cr}=\omega_{a}/2$ (details in Supplementary Information). For $\omega_{H}\leq\omega_{cr}$, the particle is synchronized with the rotating magnetic field (“phase-locked”), i.e. $\omega_{p,z}=\omega_{H}$. Otherwise, when $\omega_{H}>\omega_{cr}$, the particle undergoes back-and-forth motion (“back-and-forth”) when the maximum magnetic torque cannot balance the response drag torque: the phase lag angle $\beta$ between $\bm{n}_{xoy}$ (the projection of $\bm{n}$ on xoyxoy-plane) and $\bm{h}$ changes periodically over time [21]. These two regimes can be observed experimentally (Fig. 1f-i). In the “phase-locked” regime (Supplementary Movie 1), the particle rotates synchronously with the magnetic field, as evidenced by the normalized angular velocity, $\omega_{p,z}/\omega_{H}\approx 1$, in Fig. 1g. In the “back-and-forth” regime (Supplementary Movie 2), the particle oscillates periodically (Fig. 1i).
As illustrated in the schematic diagram of Fig. 2a, the dynamics of the particle preferred magnetization direction, $\bm{n}$, is fundamentally characterized by the phase lag angle, $\beta$, with respect to $\bm{h}$ on the xoyxoy plane. The governing equations, Eqs. (1), can be simplified and explicitly written as (derivations in Methods)
@@
Geometrically, we have $\beta=\omega_{H}t-\psi$. The phase lag angle $\beta$ represents the particle rotational displacement in the co-rotating frame of the magnetic field, while $\psi$ is the rotation angle of the projection $\bm{n}$ on the xoyxoy-plane with respect to the $x$-axis, which describes the particle rotational displacement in the laboratory frame. The turbulent vorticity signal, $\bm{\omega}_{f}$, represents the vorticity experienced by the particle center along its trajectory. This simplified representation, Eq. (2), separates different contributions and clearly shows that the rotational motion of the particle results from the interplay between the magnetic field rotational drag, the magnetic torque, and the turbulence drag torque (Fig. 2a).
@@
Now we wonder how this picture will change when turbulence is taken into consideration. The turbulence induces fluctuations on the small particles of a typical magnitude $\omega_{\eta}=1/\tau_{\eta}$, the Kolmogorov characteristic frequency. Experimentally, we turn on the turbulence and vary the intensity of the magnetic field. By statistically analyzing the probability density function (PDF) of the particle angular velocity $\bm{\omega}_{p}$, a dominant peak is observed in the $z$-direction at approximately $\omega_{p,z}\approx\omega_{H}$ (Fig. 2b-d, purple circles), while the angular velocities in the $x$ and $y$ directions are similar to each other but significantly different from that in the $z$-direction (Fig. 2b-d, yellow and green circles).
To understand the influence of turbulence on the particle dynamics further, we carry out a systematic numerical investigation. To model the motion of light particles in HIT, we first perform direct numerical simulations using a pseudo-spectral method [52, 53, 11] to obtain $\bm{\omega}_{f}$ (see Methods for details). The resulting vorticity data is then incorporated into the theoretical model, Eqs. (1), which is integrated using a third-order Runge-Kutta method.
The simulated particle rotation is validated against the experimental results (Fig. 2b-d, Supplementary Movies 3-5, respectively). Notably, the theoretical model (squares) accurately captures the essential statistical physics (peak positions) of the particle stochastic rotational dynamics in the experiments (circles). The slight discrepancies in the PDF shapes arise from polydispersity in particle shape and magnetic properties in experiments which influences how individual particles respond to an external magnetic field, whereas the simulations assume idealized spherical particles with the same magnetic characteristics for each particle (Methods and Supplementary Information). Further comparisons between simulation and experimental results all demonstrate good qualitative agreement (Supplementary Information, Supplementary Movies 9 and 10).
@@
With the theoretical model capturing the core dynamics of this complex system, and with simulations allowing exploration of a broader parameter space and finer details of particle rotation, we systematically investigate how turbulence-induced vorticity fluctuations and the deterministic magnetic forcing shape particle rotational dynamics.
Beyond the previously identified “phase-locked” and “back-and-forth” regimes, we uncover a third regime: the turbulence-dominated regime, observed at small $\omega_{a}$ w.r.t. $\omega_{\eta}$, when the effect from the magnetic field is negligible. Here, the particle rotation is primarily driven by turbulence, and its normalized angular velocity, $\omega_{p}/\omega_{H}$, fluctuates randomly around zero with no net angular drift (Fig. 3b), as evidenced by the random orientation distribution of the particle orientation, $\bm{n}$ (Fig. 3d, Supplementary Movie 6). Consequently, the phase lag $\beta$ and its derivative are also randomized (Fig. 3c). In the “phase-locked” regime ($\omega_{a}\gg\omega_{\eta}$, $\omega_{H}<\omega_{cr}$), the particle phase lag, $\beta$, remains constant, meaning $\dot{\beta}=0$ (Fig. 3g). The particle angular velocity along the magnetic field rotation axis ($z$-axis) is locked to $\omega_{H}$ (Fig. 3f, green line). The particle orientation is strongly confined, exhibiting 2D behavior (Fig. 3h, Supplementary Movie 7). In the “back-and-forth” regime ($\omega_{a}\gg\omega_{\eta}$, $\omega_{H}>\omega_{cr}$), the particle undergoes periodic acceleration and deceleration (Fig. 3j-k, inset), forming distinct looping patterns in its orientation trajectory (Fig. 3l, Supplementary Movie 8). These regimes are summarized in a phase diagram, characterized by the normalized variance of the particle angular velocity in the $z$-direction, $\langle\omega_{p,z}^{2}-\langle\omega_{p,z}\rangle^{2}\rangle/\omega_{\eta}^{2}$, where $\langle\cdot\rangle$ represents a time average. The noiseless critical frequency $\omega_{cr}=\omega_{a}/2$ is marked by a black dot-dashed line. Unlike in a quiescent fluid, the phase-locked regime (Fig. 3m, purple region) is bounded by the blue line, predicted by the scaling argument based on the simplified governing equation of Eq. (2): $\omega_{a}=2\omega_{H}+\omega_{\eta}$ (detailed reasoning can be found in Methods). This shift in the subcritical boundary, relative to the noiseless case ($\omega_{a}=2\omega_{H}$), accounts for the influence of turbulence fluctuations, which have a standard deviation on the order of $\omega_{\eta}$.
@@
To determine whether SR occurs, we focus on the $\beta$ and its behavior across different parameter settings. We summarize the time-averaged value of $\langle\dot{\beta}\rangle/\omega_{\eta}$ in one phase diagram (Fig. 4a), with $\omega_{\eta}/\omega_{a}$ (vertical axis) quantifying the relative turbulent fluctuation intensity (noise) to the applied magnetic field strength, and $\omega_{H}/\omega_{a}$ (horizontal axis) distinguishing noiseless subcritical, critical (i.e., $\omega_{H}/\omega_{a}=1/2$, black dot-dashed line), and supercritical states.
@@
For a fixed $\omega_{H}/\omega_{a}$ in the subcritical regime ($\omega_{H}/\omega_{a}<1/2$), the particle rotation in the co-rotating frame $\langle\dot{\beta}\rangle/\omega_{\eta}$ exhibits a non-monotonic dependence on noise intensity (Fig. 4b), with a pronounced resonant peak at $\omega_{\eta}/\omega_{a}\approx 1$, indicating that resonance arises at the transition from the “phase-locked” regime to the “turbulent-dynamics” regime. This is a clear suggestion that the observed resonance behavior results from the cooperation between turbulent fluctuations and magnetic forcing. Stochastic resonance is also clearly observed experimentally, as shown in Fig.4c: diamonds denote the experimental results at $\omega_{H}/\omega_{a}=0.12\pm 0.01$ (corresponding to the points marked in the phase diagram in Fig. 4a), while circles represent the simulation results at a comparable value of $\omega_{H}/\omega_{a}$ for direct comparison. The experimental and simulation results agree reasonably well; at the resonance peak, the experimental curve appears flatter which can be attributed to effects of polydispersity of the particles used in experiments. The resonance mechanism can be interpreted via an effective potential well framework (Fig. 4d), where the magnetic field induces a cosine-type potential (i.e., $\frac{\omega_{a}}{4}\cos{2\beta}$) modulated by a linear term (i.e., $\omega_{H}\beta$), while turbulence acts as a stochastic noise.
@@
For $\omega_{H}/\omega_{a}<1/2$ (Fig. 4e), the phase lag angle is initially trapped in the potential minimum at weak noise levels (small $\omega_{\eta}/\omega_{a}$), exhibiting only mild fluctuations locally. As the turbulence-induced fluctuations become comparable to the magnetic field strength ($\omega_{\eta}/\omega_{a}\approx 1$), preferential unidirectional escape occurs. This clearly identifies the observed resonant behavior as an example of stochastic resonance: a cooperative interplay between stochastic forcing and the external deterministic forcing. This phenomenon can be thought of as the particle “surfing” the rotating magnetic field, aided by the turbulent vorticity. At even higher noise levels, randomization overrides the cooperative effect, causing bidirectional escape and a decline in $\langle\dot{\beta}\rangle$.
@@
At the critical threshold $\omega_{H}/\omega_{a}=1/2$, the potential well vanishes, rendering $\beta$ critically sensitive to any perturbations (Fig. 4g). In the supercritical range ($\omega_{H}/\omega_{a}\gg 1/2$), the potential well inherently favors persistent phase slipping, regardless of noise intensity (Fig. 4h), such that the SR mechanism vanishes in this regime.